crispyrat:

1/(sqrt(36)+sqrt(27))+1/(sqrt(27)+sqrt(18))+1/(sqrt(18)+sqrt(9)) I need help, idk how to begin with this problem

AZ:

\(\dfrac{1}{\sqrt{36}+\sqrt{27}}+ \dfrac{1}{\sqrt{27}+\sqrt{18}}+\dfrac{1}{\sqrt{18}+\sqrt{9}}\) Is this your question? You're supposed to simplify this?

crispyrat:

yes

AZ:

So let's simplify some things first what is \(\sqrt{36}\) what is \(\sqrt{9}\)

crispyrat:

ok so its 6, and 3

AZ:

Good so we have \(\dfrac{1}{6+\sqrt{27}}+ \dfrac{1}{\sqrt{27}+\sqrt{18}}+\dfrac{1}{\sqrt{18}+3}\) How can we re-write \(\sqrt{27}\) and \(\sqrt{18}\) remember that 27 = 9*3 and 18 = 9*2 \(\sqrt{a\times b} = \sqrt{a}\times \sqrt{b}\) and \(\sqrt{x^2} = x\) so how can we simplify \(\sqrt{27}\) and \(\sqrt{18}\)

crispyrat:

as 3sqrt(3) and 3sqrt(2)?

AZ:

Exactly! So now we have \(\dfrac{1}{6+3\sqrt{3}}+ \dfrac{1}{3\sqrt{3}+3\sqrt{2}}+\dfrac{1}{3\sqrt{2}+3}\)

AZ:

Now the only way to add the fractions is to have the denominator And the only way to get the same denominator is by multiply both the numerator and denominator by a term so that way you can add them so for example, 1/x + 1/y you multiply 1/x by y/y you multiply 1/y by x/x and you would get x/(xy) + y/(xy) and now you can add them and get (x+y)/ (xy)

AZ:

So \(\dfrac{1}{6+3\sqrt{3}}+ \dfrac{1}{3\sqrt{3}+3\sqrt{2}}+\dfrac{1}{3\sqrt{2}+3}\) Let's add two fractions first and then simplify and then add the third one later \(\dfrac{1}{6+3\sqrt{3}}+ \dfrac{1}{3\sqrt{3}+3\sqrt{2}}\) \(=\left(\dfrac{1}{6+3\sqrt{3}}\times\dfrac{3\sqrt{3}+3\sqrt{2}}{3\sqrt{3}+3\sqrt{2}}\right) + \left(\dfrac{1}{3\sqrt{3}+3\sqrt{2}}\times\dfrac{6+3\sqrt{3}}{6+3\sqrt{3}}\right)\)

AZ:

Can you simplify it? What do you get after you multiply it and then add them

crispyrat:

so you would have (3sqrt(3)+3sqrt(2)+6+3sqrt(3))/27+18√2+18√3+9√6=(6sqrt(3)+6+3sqrt(2))/(27+18√2+18√3+9√6 )

Angle:

yes, I think that looks good so far then you want to give that mess a common denominator with the 1/(3√2 + 3) chunk

Angle:

\(\large \frac{6 \sqrt3 + 2 \sqrt2+6}{9 \sqrt6 + 18\sqrt3 + 18 \sqrt2 + 27}\times \frac{3\sqrt2+3}{3\sqrt2 +3}~~~+~~~\frac{1}{3\sqrt2 +3}\times\frac{9 \sqrt6 + 18\sqrt3 + 18 \sqrt2 + 27}{9 \sqrt6 + 18\sqrt3 + 18 \sqrt2 + 27}\)

Angle:

tbh it might make our lives easier if we factored out a 1/3 from everything

crispyrat:

oh how could i do that :p

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) Angle \(\large \frac{6 \sqrt3 + 2 \sqrt2+6}{9 \sqrt6 + 18\sqrt3 + 18 \sqrt2 + 27}\times \frac{3\sqrt2+3}{3\sqrt2 +3}~~~+~~~\frac{1}{3\sqrt2 +3}\times\frac{9 \sqrt6 + 18\sqrt3 + 18 \sqrt2 + 27}{9 \sqrt6 + 18\sqrt3 + 18 \sqrt2 + 27}\) \(\color{#0cbb34}{\text{End of Quote}}\) If we factor out a 1/3 maybe we'll get something like this \(\large \frac{1}{3}(\frac{6 \sqrt3 + 2 \sqrt2+6}{3 \sqrt6 + 6\sqrt3 + 6 \sqrt2 + 9}\times \frac{\sqrt2+1}{\sqrt2 +1}~~~+~~~\frac{1}{\sqrt2 +1}\times\frac{3 \sqrt6 + 6\sqrt3 + 6 \sqrt2 + 9}{3 \sqrt6 + 6\sqrt3 + 6 \sqrt2 + 9})\)

Angle:

there's a good chance I made a typo or something though @AZ

Angle:

yeah, definitely a typo

AZ:

In the numerator you accidentally put 2sqrt(2) instead of 3sqrt(2)

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) Angle \(\large \frac{6 \sqrt3 + 3 \sqrt2+6}{9 \sqrt6 + 18\sqrt3 + 18 \sqrt2 + 27}\times \frac{3\sqrt2+3}{3\sqrt2 +3}~~~+~~~\frac{1}{3\sqrt2 +3}\times\frac{9 \sqrt6 + 18\sqrt3 + 18 \sqrt2 + 27}{9 \sqrt6 + 18\sqrt3 + 18 \sqrt2 + 27}\) \(\color{#0cbb34}{\text{End of Quote}}\) If we factor out a 1/3 maybe we'll get something like this \(\large \frac{1}{3}(\frac{6 \sqrt3 + 3 \sqrt2+6}{3 \sqrt6 + 6\sqrt3 + 6 \sqrt2 + 9}\times \frac{\sqrt2+1}{\sqrt2 +1}~~~+~~~\frac{1}{\sqrt2 +1}\times\frac{3 \sqrt6 + 6\sqrt3 + 6 \sqrt2 + 9}{3 \sqrt6 + 6\sqrt3 + 6 \sqrt2 + 9})\)

Angle:

that makes more sense because then you can factor out a 3 in both the numerator and denominator to simplify it a bit if you wanted it or we can brute for this

Angle:

If we factor out a 1/3 maybe we'll get something like this \(\large \frac{1}{3}(\frac{2 \sqrt3 + \sqrt2+2}{\sqrt6 + 2\sqrt3 + 2 \sqrt2 + 3}\times \frac{\sqrt2+1}{\sqrt2 +1}~~~+~~~\frac{1}{\sqrt2 +1}\times\frac{\sqrt6 + 2\sqrt3 + 2 \sqrt2 + 3}{\sqrt6 + 2\sqrt3 + 2 \sqrt2 + 3})\)

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Exactly! So now we have \(\dfrac{1}{6+3\sqrt{3}}+ \dfrac{1}{3\sqrt{3}+3\sqrt{2}}+\dfrac{1}{3\sqrt{2}+3}\) \(\color{#0cbb34}{\text{End of Quote}}\) Honestly, on second thought, it would probably be so much easier if we just rationalized the fractions \( \dfrac{1}{6+3\sqrt{3}} = \dfrac{1}{6+3\sqrt{3}} \times \dfrac{6-3\sqrt{3}}{6-3\sqrt{3}} \) \( \dfrac{6-3\sqrt{3}}{(6+3\sqrt{3})(6-3\sqrt{3})} = \dfrac{(6-3\sqrt{3})}{36-27} = \dfrac{6-3\sqrt{3}}{9} \) and then similarly for the other two fractions

Angle:

sorry, we're probably confusing you

crispyrat:

wait wouldn't the one that AZ wrote be the same as before cus 6-3sqrt(3)/6sqrt(3) is the same as 1 or am i missing something

AZ:

That is the same as 1 We're multiplying by that because it's 1 and we're trying to use the property (a-b) * (a+b) = a^2 - b^2 so that way we can get rid of the square roots in the denominator and bring it to the numerator and you see that by doing that, we have a nice number in the denominator

AZ:

Found a nice image that might explain it slightly better

1 attachment
AZ:

(I assumed that they meant to write 6-3sqrt(3) but made a typo)

crispyrat:

yea

Angle:

I'm obviously not awake yet

crispyrat:

wait so if i do the samething for 1/3sqrt(2)+3sqrt(3) i would get (3sqrt(2)-3sqrt(3))/-9 I think

Angle:

yes

AZ:

yeah that could be written as (3sqrt(3) - 3sqrt(2)) / 9

AZ:

if you wanted to get rid of that negative sign

crispyrat:

so then for 1/(3sqrt(3)+3) it is (3sqrt(2)-3)/9

Angle:

yes

crispyrat:

and then i add them all together. cool. :)

Angle:

so much easier than the common denominator mess ;-;

crispyrat:

i got 3 and that over 9 is 1/3 thanks so much

AZ:

now put it all together \(\dfrac{6-3\sqrt{3}}{9} + \dfrac{3\sqrt{3} - 3\sqrt{2}}{9} + \dfrac{3\sqrt{2}-3}{9}\)

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angle so much easier than the common denominator mess ;-; \(\color{#0cbb34}{\text{End of Quote}}\) yes ;-; idk what I was thinking at first ;--;

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @crispyrat i got 3 and that over 9 is 1/3 thanks so much \(\color{#0cbb34}{\text{End of Quote}}\) Good job!